I have an equation of 4x=36 I looked at other examples of how to do it but I dont seem to understand how to figure it out, v/3 = -11 this is another one as well this is on a worksheet not a textbook
Answer by checkley71(8403)

4x=36 divide both sides by 4.
4x/4=36/4
x=36/4
x=9 answer.
------------------------
v/3=11 now cross multiply.
v=3*11
v=33 answer.